Circular No. 8951

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2671 OPHIUCHI = NOVA OPHIUCHI 2008 No. 2 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

nova reported on IAUC 8950 is being assigned the variable-star designation V2671 Oph.

K. Kinugasa, H. Taguchi and S. Honda, Gunma Astronomical Observatory (GAO); and H. Yamaoka, Kyushu University, write that a lowresolution spectrogram (range 400–800 nm; resolution ∼ 1.0 nm) of V2671 Oph was obtained with the GAO 1.5-m telescope (+ GLOWS) on June 6.6 UT. The spectrum shows a relatively red continuum and Balmer emission lines without P-Cyg profiles; also present are a strong O I emission line and many Fe II emission lines. This confirms that the object is a nova after maximum.

V2487 OPHIUCHI A. Pagnotta, B. E. Schaefer, and L. Xiao, Louisiana State University,

report their discovery of a previously unknown eruption of the nova V2487 Oph (Nova Oph 1998) in the Harvard College Observatory archival photograph collection. Plate AM505 shows V2487 Oph at B = 10.3 on 1900 June 20. This is close to its peak magnitude in 1998, which was reported to be 9.5 in the more recent outburst. They identified V2487 Oph as a strong recurrent nova candidate because of its low outburst amplitude, very fast decline rate, high expansion velocity, and the presence of high-excitation lines in its outburst spectrum. V2487 Oph was also identified as a good candidate by Hachisu et al. (2002, ASP Conf. Proc. 261, 629), based on characteristics of its light curve. With this discovery, V2487 Oph becomes the tenth known galactic recurrent nova.

COMET C/2007 W1 (BOATTINI) Total-magnitude and coma-diameter estimates (cf. IAUC 8945): May

21.37 UT, 5.8, 15′ (M. Mattiazzo, Castlemaine, Victoria, 8×40 binoculars); 23.95, 5.7, 20′ (M. Goiato, Araa̧tuba, Brazil, 7×50 binoculars); May 28.39, 5.2, 18′ (Mattiazzo); June 2.40, 4.9, 20′ (Mattiazzo); 4.92, 5.3, 15′ (W. Souza, São Paulo, Brazil, 7×50 binoculars); 5.90, 5.5, 15′ (A. Amorim, Florianopolis, Brazil, 10×50 binoculars); 6.94, 5.1, 18′ (Goiato; moonlight).

2008 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green